# “Feeling that I was Collaborating with Them:” A 20-year Systematic Literature Review of Social Virtual Reality Leveraging Collaboration

### v2

@@ 5.2. RQ2: Gaps and future directions
While the reviewed articles show the potential of Social VR for collaboration, they also expose several gaps that should be addressed to unlock its full potential. Most articles focused on short-term sessions or dyads, limiting our understanding of how immersive collaboration unfolds over extended periods or within large, diverse teams. Only a few studies examined repeated use, showing that presence can improve over time (Khojasteh and Won, 2021). Future research should explore long-term use cases and organizational contexts to better assess how Social VR supports ongoing coordination, relationship development, and task management over time.
Another major gap is the lack of standardized frameworks for evaluating collaboration in Social VR. Most studies relied predominantly on self-reported data, with limited behavioral or trace data. While CMC research has long used message logs or performance data, few articles in our review leverage these metrics. More robust, multi-modal methods are needed to capture the complexity of Social VR collaboration. Moreover, most samples were relatively small and homogeneous, limiting generalizability. Future research should examine how user diversity shapes collaboration.
Trust, while fundamental to collaboration, was notably underexplored in these articles. Only two studies in our corpus directly examined how trust develops in Social VR (Torro and Pirkkalainen, 2023; Yassien et al., 2020). These highlighted how avatar expression, proximity, and personal space influence trust. Unlike CMC platforms, where trust often develops via sustained text or asynchronous interactions, trust in VR is embodied and experiential. Understanding how team dynamics unfold in Social VR applications calls for new research on how immersive platforms can contribute to forming trust.
While many studies aimed to recreate physical-world dynamics, many articles encouraged designers to think beyond replication. As illustrated by the articles, Social VR can support forms of collaboration that transcend physical limitations, such as manipulating unreal objects, customizing users’ bodies, and using AI features and tools in immersive ways. Inspired by (Grønbæk et al., 2024; Irlitti et al., 2024), future research should study how immersive collaboration can benefit from VR’s ability to break physical restrictions: supporting asymmetric views, invisible tools, or gravity-defying interactions that foster creativity and engagement. This aligns with the “Beyond Being There” perspective (Hollan and Stornetta, 1992), which advocates for systems that not only simulate co-located interaction but also go further by introducing novel capabilities not possible in physical space.
The articles also raised concerns about usability and fatigue. Many users reported dizziness or discomfort after prolonged VR sessions. Hybrid solutions that combine immersive and non-immersive modes might offer more sustainable approaches to collaboration, especially for long-term or high-frequency use. For instance, teams might brainstorm using VR and then switch to 2D tools for task refinement. Adaptive systems that respond to user context could leverage the benefits of immersion while mitigating its limitations. Furthermore, as Social VR adoption grows, privacy and security concerns must be considered. Sensitive data during collaborations requires mechanisms that protect personal data while preserving presence and immersion.
As Social VR and Mixed Reality evolve, CSCW researchers should move beyond comparing these platforms to traditional online communication tools. Instead, future work should examine how new possibilities provided by immersive systems, such as enabling users to customize shared spaces, collaborating with embodied AI agents (Malik et al., 2020; Wang et al., 2020), and supporting new forms of teamwork that go beyond simulating in-person meetings.
@@ 5.3. Limitations and Future Work
It is important to acknowledge the limitations of this paper. First, our systematic review was scoped to three databases only. We did not include articles indexed in other systems (e.g., Scopus, Google Scholar), pre-prints (e.g., arXiv), or non-academic sources. Second, the review focused on papers from the last 20 years. Earlier work could offer additional context on the evolution of Social VR. Third, the inclusion process and thematic analysis were conducted by only one researcher, introducing bias to the selection process. We established meetings to revise and discuss the researcher’s decisions and findings. Lastly, future reviews could explore how other Extended Reality technologies have contributed to collaboration. Despite these limitations, our review advances the field by consolidating emerging work on embodied, immersive collaboration and offering a roadmap for future research in CSCW.
@@ 6. Conclusion
This systematic literature review provides a comprehensive analysis of how Social VR has been studied as a medium for collaboration. Drawing from 62 relevant articles, we identified key themes illustrating how Social VR facilitates collaboration at the individual, group, and environmental levels. The findings highlight its potential to enhance teamwork by fostering immersion, enabling expressive and customizable avatars, and facilitating social presence through nonverbal and spatial cues. Yet, important gaps remain in studying long-term collaboration dynamics, team-level processes such as trust-building, and the role of avatar design on shaping self-presentation and group cohesion.
Based on the themes and findings, we encourage researchers and practitioners to investigate sustained collaboration across diverse teams, develop standardized evaluation frameworks, and embrace the unique affordances of Social VR to create more engaging, inclusive, and transformative collaborative experiences.

### v3

@@ 5.2. RQ2: Gaps and Future Directions
The corpus also reveals theoretical and methodological gaps in Social VR research (RQ2). First, most empirical studies involved dyads and small groups in short periods. While these studies have shed light on presence, trust, and coordination, they raise questions about scalability and long-term effects. How can Social VR support larger teams, collectives, or organizational structures? How should multi-user work be managed and distributed? How persistent can these collaborations be after days, weeks, or months? CSCW research has long examined how group size alters coordination, governance, and knowledge sharing (Harris et al., 2019). Yet, these insights are largely absent from current Social VR research. More empirical evidence is needed, including longitudinal studies, field deployments, and diverse samples. In particular, characterizing VR’s distinct affordances—such as spatial partitioning, altered scales and distances, and subgroup interactions—can provide new possibilities to enable collectives that are constrained by physical spaces and require embodied interactions.
Second, most studies have relied on convenience samples—often college students and Western contexts—which limits the generalizability of their findings. This raises broader concerns that many Social VR systems may be implicitly designed around Western norms and expectations, potentially overlooking cultural differences in interaction styles, privacy preferences, or embodiment needs that could influence how people from diverse backgrounds experience and benefit from Social VR. Future research should recruit broader participant pools, including non-student populations, individuals with limited exposure to VR technologies, marginalized groups, and participants from varied cultural contexts. Broadening these samples will help diversify and strengthen the external validity of Social VR research.
Third, the review also highlights the lack of a robust ecosystem of applications in Social VR specifically designed for collaboration. Most studies relied on general-purpose platforms (e.g., AltspaceVR, VRChat), and some researchers tested them for specific collaborative purposes. While there is variety in productivity applications (e.g., Slack, MS Teams, Gmail), Social VR lacks dedicated, interoperable tools designed to integrate into everyday collaborative workflows. As suggested by our themes, Social VR requires not only designing the productivity features and tools to enable work but also considering how users experience collaboration contingent on individual, team, and environmental factors.
The corpus analysis also revealed that ethics and accessibility require consistent analysis, reporting, and discussion. HMD-based systems present significant barriers for users with visual, auditory, or motor impairments, raising concerns about adaptability and inclusion. Moreover, more discussions are needed to address ethical challenges, such as surveillance, harassment, and biometric data collection (e.g., motion, eye data). Many of these papers rarely specified what signals were recorded, for how long they were retained, what protections were in place, or whether accommodation features were available. Future studies should make these design choices, data collection practices, and accessibility features explicit. Accessibility is essential to prevent Social VR from becoming exclusionary.
These mixed findings raise the broader question of “why Social VR?” That is, under what conditions does it provide advantages over video calls or in-person meetings, and when does it fall short? Based on this corpus, the clearest advantages emerge in tasks that require room-scale layout, shared 3D manipulation, joint navigation, or view sharing—capabilities beyond standard video conferencing. Many studies compared Social VR to video calls, with few comparing to in-person collaboration. As such, the current research leaves limited evidence about how these modes differ and how Social VR can reconfigure face-to-face dynamics. Overall, the studies indicate that Social VR is most effective when the collaborative task depends on embodied and spatial interaction. However, many applications continue to attempt to directly mimic physical offices or simply enable conversational dynamics, leading to increased frustration and costs compared to existing solutions.
5.3. Design Implications
Together, these themes reveal design conditions shaping Social VR collaboration: how expressive embodied interaction shapes individual experience, how groups coordinate and build trust, and how environments support joint action. Building on these thematic insights, we identify three design implications that translate these patterns into actionable directions for future systems.
Prioritizing Interaction Fidelity over Visual Fidelity
Across the corpus, breakdowns in presence and social presence are frequently caused by low-fidelity motion, delayed or unnatural hand alignment, missing or awkward expressive cues, and limited support for natural gestures. These shortcomings hindered both the sense of “being there” and of “being there with others,” reducing engagement and limiting the authenticity of collaborative exchanges described in the sub-themes. Improving core interaction fidelity—e.g., stable object manipulation, predictable locomotion, low-latency input, local interpolation of movement, and expressive baseline cues when eye or face tracking is unavailable—should be a priority rather than just making the avatars more “realistic.” Collaboration benefits more from responsive embodied action than from avatar photorealism. Future systems should invest in reliable and expressive interaction as the foundation for collaborative behaviors in Social VR.
Centering Collaboration around Shared Objects, Tools, and Spatial Coordination Mechanisms
Across the studies, collaboration improved when groups could easily create, manipulate, and reference shared artifacts. When tools were rigid, limited, or difficult to access, teams reported reduced cohesion, situational awareness, and engagement. The themes demonstrated how virtual objects serve not only as functional resources but also as social anchors, enabling collaborators to maintain a mutual focus and coordinate their actions. Additionally, spatial affordances—such as view-sharing, group-preserving teleportation, and visibility-balanced layouts—played a major role in enabling common ground. Thus, Social VR systems should emphasize robust, low-friction multi-user tools, persistent artifacts, and spatial mechanisms that enhance group awareness. These features support collaborative fluency and reduce “re-entry costs” that appear when teams must re-establish context.
Designing for Trust, Safety, Accessibility, and Transparency
Findings across themes highlight that trust, safety, and comfort strongly influence participation and sustained collaboration. Privacy violations, proximity discomfort, unpredictable avatar movements, and inaccessible interaction modes eroded trust and undermined users’ sense of social ease. Similarly, missing accommodations—such as captions, alternative input options, or personal-space controls—limited inclusivity and led to uneven participation. A design implication is that collaborative effectiveness depends on whether users feel protected, respected, and able to interact on their own terms. Incorporating accommodation elements should be a design priority in future systems. Moreover, many articles from the corpus provided limited detail about behavioral signal collection and data handling. Transparent reporting to the users of what signals are captured, how long they are retained, and what protections are applied can improve trust. Embedding safety, accessibility, and transparency directly into system design will broaden participation and help ensure Social VR remains comfortable and sustainable.
@@ 5.4. Limitations and Future Work
It is important to acknowledge the limitations of this paper. First, we focused on peer-reviewed venues in the ACM Digital Library, IEEE Xplore, and Web of Science, core outlets for CSCW, HCI, and Social VR research. This ensured a reliable corpus but excluded pre-prints, dissertations, reports, and other gray literature. databases (e.g., Scopus, arXiv) may provide wider coverage. Future reviews could include non-peer-reviewed work—with explicit screening procedures—as industry-influences VR development.
Second, we limited our corpus to the period from 2004 to 2024 to capture the era of HMD-based Social VR. Earlier research on CVEs, CAVEs, and other screen-based platforms were excluded. While this decision allowed us to concentrate on the recent technologies shaping the current Social VR ecosystem, future reviews could revisit how pre-2004 work informs current systems and identify which insights have been lost or should be re-studied.
Third, our focus was on Social VR, defined as multi-user experiences mediated through HMDs. Other immersive technologies, such as Augmented Reality (AR) or Mixed Reality (MR), were excluded from our review. This maintained conceptual clarity in treating VR as a distinct medium. However, this restricts generalizability to new Extended Reality (XR) modalities. As the XR ecosystem becomes more widespread, future work should investigate how this broad spectrum of immersive technologies supports collaboration, potentially identifying new affordances and frameworks.
Fourth, our reliance on collaboration and work-related search terms may have excluded studies on Social VR in specific contexts, such as with older adults or marginalized groups, but did not explicitly frame these as “collaboration.” Future reviews should broaden collaboration definitions to include contexts where individuals pursue shared goals.
Fifth, scoping reviews are designed to map breadth rather than evaluate evidence strength. Because the focus is on coverage rather than depth, the synthesis remains descriptive, identifying patterns across studies without assessing reliability of individual findings or comparing effect sizes. Future systematic reviews or meta-analyses could build on this to rigorously assess evidence quality and consistency.
Lastly, screening and thematic coding were led by one researcher, with decisions discussed in meetings. While having meetings increased the consistency and robustness of these themes, the emerging codes depended on the first researcher, introducing bias in the analysis. Future studies should adopt multi-coder approaches or other assisted tools to prevent bias and avoid missing relevant codes.
@@ 6. Conclusion
This scoping review provides a comprehensive analysis of how Social VR has been studied as a medium for collaboration. Drawing from 62 relevant articles, we identified three conceptual levels (i.e., individual, team, and affordances) that shape how collaboration unfolds in Social VR. Our findings highlight its potential to enhance teamwork by fostering presence, supporting expressive and customizable avatars, and enabling social presence through nonverbal and spatial cues. The review also revealed important gaps in this research area, such as examining long-term collaboration dynamics, team-level processes such as trust-building, and the influence of avatar design on shaping self-presentation and group identity.
Building on these insights, this paper encourages researchers and practitioners to move beyond short-term experimental designs and explore how Social VR can support sustained collaboration across diverse teams, domains, and cultural contexts. Taken together, this review highlights both the promise and the limitations of Social VR as a medium for facilitating collaboration. For CSCW, it highlights a renewed opportunity to shape the theories, methods, and design practices that will determine whether Social VR becomes a fleeting novelty or a sustainable foundation for the future of remote collaboration.
Appendix A List of included articles in the final corpus
Overview of the studies included in the final corpus, along with their corresponding themes and subthemes.
Table A1. Overview of the final Social VR collaboration corpus.
Long table listing all articles included in the final Social Virtual Reality collaboration corpus. Each row reports an article identifier, title, reference citation, and the sub-themes to which the article was assigned.
ID Title Ref Themes
1 Conveying perspective in multi-user virtual reality collaborations (Freiwald et al., 2020) Realistic & Functional Environments
2 CakeVR: A Social Virtual Reality (VR) Tool for Co-designing Cakes (Mei et al., 2021) Sense of Presence; Social Presence
3 Falling Asleep Together: What Makes Activities in Social Virtual Reality Meaningful to Users (Maloney and Freeman, 2020) Real Collaboration
4 Enhanced Auditoriums for Attending Talks in Social Virtual Reality (Weissker et al., 2023) Social Presence
5 Just undo it: exploring undo mechanics in multi-user virtual reality (Rasch et al., 2024) Virtual Tools & Objects
6 Remote and Collaborative Virtual Reality Experiments via Social VR Platforms (Saffo et al., 2021) Realistic & Functional Environments
7 Communication in Immersive Social Virtual Reality: A Systematic Review of 10 Years’ Studies (Wei et al., 2022) Social Presence
8 Working Together Apart through Embodiment: Engaging in Everyday Collaborative Activities in Social Virtual Reality (Freeman et al., 2022) Realistic & Functional Environments
9 Give-Me-A-Hand: The Effect of Partner’s Gender on Collaboration Quality in Virtual Reality (Yassien et al., 2021) Avatar Design
10 Enhancing collaborative manipulation through the use of feedback and awareness in CVEs (García et al., 2008) Virtual Tools & Objects
11 A collaborative VR visualization environment for offshore engineering projects (dos Santos et al., 2011) Virtual Tools & Objects
12 Effect of gender on immersion in collaborative iODV applications (Kallioniemi et al., 2017) Avatar Design
13 A review on effective closely-coupled collaboration using immersive CVE’s (Otto et al., 2006) Virtual Tools & Objects
14 Subjective evaluation of group user QoE in collaborative virtual environment (CVE) (Moharana et al., 2022) Social Presence
15 Cone of Vision as a Behavioural Cue for VR Collaboration (Bovo et al., 2022) Social Presence
16 Inter-brain Synchrony and Eye Gaze Direction During Collaboration in VR (Gumilar et al., 2022) Social Engagement
17 Impact of Augmented Engagement Model for Collaborative Avatars on a Collaborative Task in Virtual Reality (Le Tarnec et al., 2024) Social Engagement
18 Visual deictic reference in a collaborative virtual environment (Duchowski et al., 2004) Social Engagement
19 A Design Space for Social Presence in VR (Yassien et al., 2020) Avatar Design; Teammate Trust
20 VRGit: A Version Control System for Collaborative Content Creation in Virtual Reality (Zhang et al., 2023) Realistic & Functional Environments
21 blue-c API: a multimedia and 3D video enhanced toolkit for collaborative VR and telepresence (Naef et al., 2004) Spatial Audio & Movement
22 Let’s Face It: Influence of Facial Expressions on Social Presence in Collaborative Virtual Reality (Kimmel et al., 2023) Social Presence
23 Identifying Cognitive and Creative Support Needs for Remote Scientific Collaboration using VR: Practices, Affordances, and Design Implications (Olaosebikan et al., 2022) Sense of Presence; Social Presence
24 Towards Immersive Collaborative Sensemaking (Yang et al., 2022b) Social Engagement
25 Designing Virtual Environments for Social Engagement in Older Adults: A Qualitative Multi-site Study (Xu et al., 2023) Social Engagement
26 Users’ locomotor behavior in collaborative virtual reality (Ríos et al., 2018) Spatial Audio & Movement
27 Invoke: A Collaborative Virtual Reality Tool for Spatial Audio Production Using Voice-Based Trajectory Sketching (Deacon and Barthet, 2023) Spatial Audio & Movement
28 LeMo: Exploring Virtual Space for Collaborative Creativity (Men and Bryan-Kinns, 2019) Realistic & Functional Environments
29 Being Social in VR Meetings: A Landscape Analysis of Current Tools (Osborne et al., 2023) Avatar Design; Real Collaboration; Realistic & Functional Environments; Virtual Tools & Objects
30 AMMP-Vis: a collaborative virtual environment for molecular modeling (Chastine et al., 2005) Social Presence
31 CollaVR: Collaborative In-Headset Review for VR Video (Nguyen et al., 2017) Virtual Tools & Objects
32 Synchrony within Triads using Virtual Reality (Miller et al., 2021) Sense of Presence
33 Sushi with Einstein: Enhancing Hybrid Live Events with LLM-Based Virtual Humans (Shoa et al., 2023) Avatar Design
34 Social Simon Effect in Virtual Reality: Investigating the Impact of Co-actor Avatar’s Visual Representation (Li et al., 2024) Avatar Design
35 Supporting social human communication between distributed walk-in displays (Roberts et al., 2004) Virtual Tools & Objects
36 Going, Going, Gone: Exploring Intention Communication for Multi-User Locomotion in Virtual Reality (Rasch et al., 2023) Spatial Audio & Movement
37 Effects of Avatar’s Symbolic Gesture in Virtual Reality Brainstorming (Ide et al., 2020) Social Presence
38 Proof-of-Concept Study to Evaluate the Impact of Spatial Audio on Social Presence and User Behavior in Multi-Modal VR Communication (Immohr et al., 2023) Spatial Audio & Movement
39 The Effects of Self- and External Perception of Avatars on Cognitive Task Performance in Virtual Reality (Kocur et al., 2020) Avatar Design
40 Dyadic interactions with avatars in immersive virtual environments: high fiving (Young et al., 2015) Avatar Design
41 Lessons From Working in the Metaverse: Challenges, Choices, and Implications from a Case Study (Park et al., 2024) Real Collaboration
42 Design principles for social exchange in social virtual reality-enabled virtual teams (Torro and Pirkkalainen, 2023) Avatar Design; Real Collaboration; Teammate Trust
43 Studying the effect of symmetry in team structures on collaborative tasks in virtual reality (Agnès et al., 2023) Social Presence
44 Working Together on Diverse Tasks: A Longitudinal Study on Individual Workload, Presence and Emotional Recognition in Collaborative Virtual Environments (Khojasteh and Won, 2021) Sense of Presence
45 Communicating Information in Virtual Reality: Objectively Measuring Team Performance (Moore et al., 2019) Social Engagement
46 Getting There Together: Group Navigation in Distributed Virtual Environments (Weissker, 2020) Spatial Audio & Movement
47 Supporting a closely coupled task between a distributed team: Using immersive virtual reality technology (Roberts et al., 2005) Virtual Tools & Objects
48 Group Navigation for Guided Tours in Distributed Virtual Environments (Weissker and Froehlich, 2021) Realistic & Functional Environments
49 Collaborative virtual reality platform for visualizing space data and mission planning (García et al., 2019) Virtual Tools & Objects
50 See-through techniques for referential awareness in collaborative virtual reality (Argelaguet et al., 2011) Virtual Tools & Objects
51 Measuring the collaboration degree in immersive 3D collaborative virtual environments (Sarmiento et al., 2014) Spatial Audio & Movement
52 Multi-Ray Jumping: Comprehensible Group Navigation for Collocated Users in Immersive Virtual Reality (Weissker et al., 2019) Spatial Audio & Movement
53 Multiparty Holomeetings: Toward a New Era of Low-Cost Volumetric Holographic Meetings in Virtual Reality (Langa et al., 2022) Realistic & Functional Environments
54 A Longitudinal Study of Task Performance, Head Movements, Subjective Report, Simulator Sickness, and Transformed Social Interaction in Collaborative Virtual Environments (Bailenson and Yee, 2006) Real Collaboration
55 Evaluating the Effect of Binaural Auralization on Audiovisual Plausibility and Communication Behavior in Virtual Reality (Immohr et al., 2024b) Spatial Audio & Movement
56 Physiological Synchrony in a Collaborative Virtual Reality Task (Moharana et al., 2023) Social Engagement
57 Subjective Evaluation of the Impact of Spatial Audio on Triadic Communication in Virtual Reality (Immohr et al., 2024a) Spatial Audio & Movement
58 Kinetic Connections: Exploring the Impact of Realistic Body Movements on Social Presence in Collaborative Virtual Reality (Kimmel et al., 2024) Avatar Design
59 Hands or Controllers? How Input Devices and Audio Impact Collaborative Virtual Reality (Adkins et al., 2024) Social Presence
60 GazeMolVR: Sharing Eye-Gaze Cues in a Collaborative VR Environment for Molecular Visualization (Darbar et al., 2024) Social Presence
61 Social virtual reality: systematic review of virtual teamwork with head-mounted displays (Dey et al., 2024) Social Presence; Real Collaboration; Virtual Tools & Objects
62 Exploring sex differences in collaborative virtual environments for participation equality and user experience (Yang et al., 2024) Avatar Design